Please help:
Tilly is graduating this spring and wishes to endow NYU with 200,000 in 14 Years. find the amount she would have to invest on graduation at 12% compound continuously to achieve this.
Thank you for taking the time to help.

Answer by nerdybill(7384)

Tilly is graduating this spring and wishes to endow NYU with 200,000 in 14 Years. find the amount she would have to invest on graduation at 12% compound continuously to achieve this.
.
Continuous Compound Interest Formula
A = Pe^(rt)
where,
P = principal amount (initial investment) (???)
r = annual interest rate (as a decimal) (.12)
t = number of years (14)
A = amount after time t (200000)
.
200000 = Pe^(.12*14)
200000 = Pe^1.68
200000/e^1.68 = P
200000/5.365555971 = P
$37274.80 = P